OpenStudy (anonymous):

inverse of F f(x) = 2 + ln x

OpenStudy (anonymous):

y = 2 + lnx inverse switch x and y and solve for y so x = 2 + lny lny=x-2 e^(lny)=e^(x-2) y=e^(x-2)

OpenStudy (anonymous):

if you don't understand that, the rule is: e^(lna)=a.. or in your case y